How do I calculate UCL/LCL (95%) from the following available data; Coefficient, Standard Error, T-Stat, P-Value.

Regression Statistics Multiple R R Square Adjusted R Square Standard Error Observations 0.9613 0.9241 0.9222 1425.3397 200 ANOVA Significance df MS Regression Residual Total 479410417802.47 95882083560.49 472.4892 39368362197.53 518778780000.00 0.00 194 199 202929702.05 Upper 95% CoefficientsStandard Erro t Stat P-value Lower 95% 45482.366 10383.543 11.088 738.388 0.014 2.546 19403.8863 3153.7202 10.4859 175.8223 0.0023 1.2209 2.340.0201 0.0012 0.2916 0.0000 0.0000 0.0383 Intercept V1 v2 3.29 1.06 4.20 6.37 2.09 V4 V5
Answer #1

df for Residual is 194

Critical t value for 95% confidence interval and df = 194 is  1.972

Upper 95% level = Coefficient + 1.972 * Standard Error

Lower 95% level = Coefficient - 1.972 * Standard Error

For Intercept,

Upper 95% level =  45482.366 + 1.972 * 19403.8863 = 83746.83

Lower 95% level = 45482.366 - 1.972 * 19403.8863 = 7217.902

For v1,

Upper 95% level = -10383.543 + 1.972 * 3153.7202 = -4164.407

Lower 95% level = -10383.543 - 1.972 * 3153.7202 = -16602.68

For v2,

Upper 95% level = 11.088 + 1.972 * 10.4859 = 31.76619

Lower 95% level = 11.088 - 1.972 * 10.4859 = -9.590195

For v3,

Upper 95% level = 783.388 + 1.972 * 175.8223 = 1130.11

Lower 95% level = 783.388 - 1.972 * 175.8223 = 436.6664

For v4,

Upper 95% level = 0.014 + 1.972 * 0.0023 = 0.0185356

Lower 95% level = 0.014 - 1.972 * 0.0023 = 0.0094644

For v5,

Upper 95% level = -2.546 + 1.972 * 1.2209 = -0.1383852

Lower 95% level = -2.546 - 1.972 * 1.2209 = -4.953615